OpenStudy (anonymous):

If (2a+3)/5 = a/2, then a+3 = ?

OpenStudy (anonymous):

The study guide gave me the answer, I just don't understand how they got 9. I am still a little confused.

OpenStudy (anonymous):

\[\frac{(2a+3)}{5} = \frac{a}{2} \rightarrow 2(2a+3) = 5a \rightarrow 4a+6=5a\] i.e.6=5a -4a i.e. 6=a or a=6 since we need to find a+3 therefore a+3= 6+3=9 i.e. a+3=9 which is the required ans.

OpenStudy (anonymous):

do u still have some problem?

OpenStudy (anonymous):

dpasingh, thanks you for your help. Just one question, how did you get a to equal 6?

OpenStudy (anonymous):

The rest makes sense to me...

OpenStudy (chmvijay):